OpenStudy (anonymous):

Radical Expressions and Equations Help: (8+√11)(8 -√11) Straight up answer if you can. Answers ------- 53 75 + 16 √11 -57 64 + √11

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow (a + b)(a - b) = a^2 - b^2 \Longleftrightarrow (8 + \sqrt{11})(8 - \sqrt{11}) = 64 - 11 = 53 }\)

Parth (parthkohli):

Oops, let me fit in the next line here: \(\Large \color{MidnightBlue}{\Rightarrow = 64 - 11 }\) \(\Large \color{MidnightBlue}{\Rightarrow = 53 }\)

OpenStudy (anonymous):

Thanks, and one more question, what math grade level is this stuff in?

Parth (parthkohli):

Idk, these are radicals. This is pre-algebra stuff I guess.

OpenStudy (anonymous):

Cool. thanks again. Closing. Might have more. QQ....

Parth (parthkohli):

Sure, turn up if you need.